if a man clear a plot of land at 3 hours, another man clear a plot of land at 5 hours and another man clear a plot of land at 7 hours, how many hours will it take 3 of the men to clear one plot of land
Answer by TimothyLamb(4379)

r = w/t
rates are additive:
1/3 + 1/5 + 1/7 = 0.6761904762
---
t = w/r
---
t = 1 / 0.6761904762
t = 1.47887323944 hours
t = 1 hour 0.47887323944 * 60 minutes
t = 1 hour 28.7 minutes
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php